OpenStudy (anonymous):

The length of an arc defined by a 50o central angle in a circle of radius 3 in. is __?__ in. (Round to the nearest hundredth).

OpenStudy (mertsj):

This is the same problem I just showed you how to do. \[\frac{50}{360}\times2\times \pi \times r = arc length\] Replace r with 3 and solve.